40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Feb 25 15:30:29 2013, yuta, Update, Alignment, Y arm locked, both colors 
    Reply  Mon Feb 25 17:58:28 2013, rana, Update, Alignment, Y arm locked, both colors 
       Reply  Mon Feb 25 22:42:32 2013, yuta, Summary, Alignment, current IFO situation IPANGdrift.png
    Reply  Mon Feb 25 21:25:14 2013, yuta, Update, Alignment, both arms locked in IR QUAD1_1045890588.png
       Reply  Tue Feb 26 01:49:27 2013, Jenne, Update, Locking, PRMI locked 
          Reply  Tue Feb 26 02:04:51 2013, Koji, Update, Locking, PRMI locked 
             Reply  Tue Feb 26 11:55:11 2013, Manasa, Update, Locking, FPMI locked 
                Reply  Tue Feb 26 15:32:47 2013, Jenne, Update, Locking, FPMI locked 
             Reply  Wed Feb 27 01:09:53 2013, yuta, Update, Locking, PRMI sideband/carrier locked 6x
                Reply  Wed Feb 27 02:44:39 2013, Jenne, Update, Locking, DRFPMI flashes 
                   Reply  Wed Feb 27 20:10:43 2013, rana, Update, Locking, DRFPMI flashes 
Message ID: 8177     Entry time: Wed Feb 27 01:09:53 2013     In reply to: 8166     Reply to this: 8179
Author: yuta 
Type: Update 
Category: Locking 
Subject: PRMI sideband/carrier locked 

[Jenne,Yuta]

We succeeded in locking PRMI in sideband and carrier.
Measured power recycling gain was ~60 power recycling gain was 4 (edited by YM on Feb 27; see elog #6947), but we have many things we cannot explain.

Snapshots:
  Here you are, Jamie.

PRMI sideband locked
POP_1045983973.bmp REFL_1045983963.bmp AS_1045983954.bmp


PRMI carrier locked
POP_1045986718.bmp REFL_1045986730.bmp AS_1045986710.bmp

  I centered POP camera and put attenuator to take these snapshots.

  Compare with previous ones.

Aug 17, 2012 elog #7213
Jun 28, 2012 elog #6886
Mar 15, 2012 elog #6421


Locking details:
 == MI only ==
  MICH: AS55_Q_ERR, AS55_PHASE_R = -12 deg, MICH_GAIN = -7, feedback to BS

 == PRMI sideband ==
  MICH: AS55_Q_ERR, AS55_PHASE_R = 24.5 deg,  MICH_GAIN = -0.05 (acquisition) -> -5 (UGF ~100 Hz), feedback to BS
  PRCL: REFL33_I_ERR, REFL33_PHASE_R = -22.65 deg, PRCL_GAIN = 4 (UGF ~120 Hz), feedback to PRM

 == PRMI carrier ==
  MICH: AS55_Q_ERR, AS55_PHASE_R = 24.5 deg,  MICH_GAIN = -0.08 (couldn't measure UGF), feedback to BS
  PRCL: REFL33_I_ERR, REFL33_PHASE_R = -22.65 deg, PRCL_GAIN = -0.3 (couldn't measure UGF), feedback to PRM


Power recycling gain:
  POPDC was 32 when PRM is misaligned, 25 when PRMI sideband locked, ~2000 when PRMI carrier locked.
  This means, power recycling gain is ~60 power recycling gain is ~4 (=POPlocked/POPmis*T_PRM=2000/30*0.06). Expected power recycling gain for PRMI is ~45, when there's no loss (see elog #6947).

  I reduced POPDC PD gain so that it doesn't saturate.


Issues:
  - We optimized AS55_PHASE_R to -12 deg by looking at MI signal. But somehow, -12 deg didn't work for PRMI.
  - Somehow, REFL11_I didn't work to lock PRCL.
  - REFL11_Q didn't work to lock MI. SNR not enough?
  - We saw POPDC flashing up to ~15000. What is this?
  - Carrier lock was not stable, we couldn't hold for more than ~30 sec. It looks like PRM moves too much when PRMI is locked.
  - Input pointing is drifting a lot in pitch. I had to re-align TT2/TT1 to the arms every ~1 hour to get good MI alignment. When I tweak TT2/TT1, both TRY and TRX gets better. I think this shows that input pointing is drifting, not the arms.


Next:
 - redo PRM/BS coil balancing
 - optimize REFL33 rotation phase
 - stabilize carrier lock somehow
 - measure PRC g-factor

ELOG V3.1.3-